Find the time per unit (in minutes) required to produce the first unit, if the slope parameter of the learning rate is 86 % and the time to complete unit 65 takes 9 minutes. (Enter your answer as a number without the units.)
Answer #1

Learning Curve formula can be stated as below

Y=a*X^b

Y=Cumulative average time to produce per unit=9 minutes

a=Time taken to produce first unit

X=number of units produced

b=Log(Learning rate)Log2=Log(0.86)/Log2=-0.217

therefore equation becomes

9=a*(65)^(-0.217)

9*65^(0,217)=a=22.266 minutes

Hence time required to produce very first unit is 22.266 units

hence the answer is 22.27